OpenStudy (peterg2001):

help please Based on the lesson, identify the first operation needed to solve the following equation. 2x + 3 = 11 addition subtraction multiplication division

OpenStudy (compassionate):

You would need to subtract 3, then divide by 2.

TheSmartOne (thesmartone):

^^^

OpenStudy (anonymous):

Make sense?

OpenStudy (peterg2001):

nr

OpenStudy (anonymous):

I find it to be kinda tricky cause you could technically divide first, but would be a pain...

OpenStudy (anonymous):

Whatever you do to one side of an equation you have to also do to the other, basics of algebra.

OpenStudy (anonymous):

We are going to try to get x by itself, so we can solve for it.

OpenStudy (anonymous):

So if you have 2x+3=11 The first thing I would do is just subtract 3 from both sides so then we would have 2x = 8 which is alot closer to having x by itself.

OpenStudy (anonymous):

At that point we could easily divide both sides by 2. 2x/2 = 8/2 therefore x = 4

OpenStudy (anonymous):

How is that? I always like to ad or subtract numbers first that are not being multiplied by variable. After variable is isolated (possibly with a multiplier out front like 2x) we can then divide to solve for x.

OpenStudy (anonymous):

Any more clear now?

OpenStudy (anonymous):

If it doesn't make sense try this video (; https://www.youtube.com/watch?v=9Ek61w1LxSc